Feder Observatory

A Blog for Minnesota State University Moorhead's Paul P. Feder Observatory

October 23, 2014 Observing Log 0

Posted on October 24, 2014 by Juan

Lauren U., Erin A., Brandon C., Brian T., and Eli H. went out to the telescope in hopes of observing an exo-planet transit, however, time ran short. So the M-29 cluster was observed instead. Skies were clear and the CCD camera actually worked. However, the computer froze when trying to get flats, darks, and biases, and had to be rebooted.

Observing Log 10/23/2014 0

Posted on October 23, 2014 by Matt

Lauren U., Erin A., Brandon C., Brian T., and Eli H. went out to the telescope in hopes of observing an exo-planet transit, however, time ran short. So the M-29 cluster was observed instead. Skies were clear and the CCD camera actually worked. However, the computer froze when trying to get flats, darks, and biases, and had to be rebooted.

Observing Log 10/21/2014 0

Posted on October 22, 2014 by Juan

Elias Holte, Erin Aadland, Lauren Ugelstad, Brandon Crowe, Brian Ternes, and Dr Craig spent around four hours behind the controls tonight with a temperature of 59.7F, focus set at 2842, and with a FWHM of 5.27. Starting around 7:45 everyone but Dr Craig (who arrived around 8:30) worked on getting the observatory up and running. After nearly an hour of troubleshooting Dr Craig came to the rescue and flipped the driver toggle switch to the “on” position. This was much appreciated.

We were able to take data on two objects tonight:

  • M29
  • M15

NOTE: Always make sure anyone in the observatory knows when you’re moving the telescope. Or wear hard hats.

Observing Log 10/21/2014 0

Posted on October 21, 2014 by Matt

Elias Holte, Erin Aadland, Lauren Ugelstad, Brandon Crowe, Brian Ternes, and Dr Craig spent around four hours behind the controls tonight with a temperature of 59.7F, focus set at 2842, and with a FWHM of 5.27. Starting around 7:45 everyone but Dr Craig (who arrived around 8:30) worked on getting the observatory up and running. After nearly an hour of troubleshooting Dr Craig came to the rescue and flipped the driver toggle switch to the “on” position. This was much appreciated.
We were able to take data on two objects tonight:
M29
M15
NOTE: Always make sure anyone in the observatory knows when you’re moving the telescope. Or wear hard hats.

Non-observing Log, 10/20/2014 0

Posted on October 20, 2014 by Juan

[Posted by Juan on behalf of the Astro 366 Students]

Erin A., Brandon C., Brian T., Eli H., and Lauren U. had a very interesting night.  The night was clear and we were ready to observe the sky, but unfortunately we ran into a problem connecting to the camera.  When we tried to connect, we received a pop up window that said “error opening camera, could not initialize the ccd camera(5).”  We managed to get both Juan Cabanela and Matt Craig on the phone at the same time.  After putting all seven of our heads together, we tried a variety of things and after being unsuccessful we gave up for the night.

Observing Log 10/19/2014 0

Posted on October 19, 2014 by Matt

Erin A., Brandon C., Brian T., Eli H., and Lauren U. had a very interesting night.  The night was clear and we were ready to observe the sky, but unfortunately we ran into a problem connecting to the camera.  When we tried to connect, we received a pop up window that said “error opening camera, could not initialize the ccd camera(5).”  We managed to get both Juan Cabanela and Matt Craig on the phone at the same time.  After putting all seven of our heads together, we tried a variety of things and after being unsuccessful we gave up for the night.

Observing Log, 10/9/2014 0

Posted on October 09, 2014 by Matt

Matt Craig, Erin Aadland, Brian Ternes and Brandon Crowe went observing tonight. Very clear night, shot several calibration shots, then science images of:

  • BU Cam (R filter)
  • Kelt-1 b (R filter)
  • NGC 1039 (BVRI)

Air temperature was 36F, focus set to 2830 and FWM around 5. We ended up just taking one image of each science object because people needed to get back to homework.

We also took out one of the 6-inch celestrons and looked at the Moon, Mizar/Alcor, Albireo, the Ring Nebula and Uranus. Technically we probably also had some photons from Pluto hit our eyes too but we couldn’t actually see it.

Observing log, 10/8/2014 0

Posted on October 09, 2014 by Matt

Nathan Walker, Elias Holte and I observed this evening. The science targets were NGC 6946, TReS-2 b and EY UMa. Seeing was very good (FWHM of 5-6 with focus of 2831, air temperature of 28F) but we hit a number of issues:

  • We forgot to turn on TCS after shooting calibration frames (got some nice star trails, though!)
  • The V filter looks very “foggy”, as it has on occasion in the past. Should probably be cleaned again.
  • Windows had to reinstall the camera driver. Again.
  • I left briefly, and when I returned the telescope was “trailing” again. Wre weren’t sure why, but resetting TCS to user default settings seemed to fix it, for a while, until…
  • …star trails, again. After trying several things I went out to the dome and noticed the hand paddle that hangs from the back of the telescope was pressing against the camera because of the orientation of the telescope. It was pushing down the “N” button (which explained the green “N” I was seeing in TCS) so I rotated the paddle 180 degrees and was back in business.
  • But only for a little while, because the camera died suddenly. Made another trip to the dome and found that the power brick had fallen of the pedestal, unplugging the camera. Plugged it back in and the rest of the night was uneventful.

Nathan and Elias left around 11:30PM, I stayed until around 7AM.

I changed the focus to 2821 around 2AM, temperature had dropped to around 20F.

 

Matt Craig

Observing Log 9/18/2014 0

Posted on September 18, 2014 by Juan

Juan and Matt out here with Lauren U, Brian T., Erin A., and Brandon C. for telescope training.  “Interesting night” in the cursed sort of “interesting” way.

Computer would not power on: Initially the computer would not power on.  We played with the cords and so on and came to the conclusion it could be either the power supply or the (watch) battery for the BIOS.  However, as the I (Juan) took the kiddos out to the dome to learn how to open up the dome, Matt got the computer to boot.  No magic.  He unplugged it for a few seconds, then plugged it in.  This had already been attempted, but hey, “interesting”…

Camera Drivers self-immolated again: Then we discovered that the computer was not recognizing the camera.  It turned out we had to install the Apogee camera driver (again, this happened a few weeks ago).  Something is seemingly wiping out the driver.

Camera focus was far off the model we have been using in the past: We next pointed the telescope using http://nova.astrometry.net/ and discovered we had the focus really far off (preventing a astrometric solution).

For a temperature of 72F we discovered we had to get the focus to 2865 (about 20 above where our previous solutions were) to get things in focus, but it was astounding seeing (FWHM of 3 pixels!).

Finally, the kiddos take pictures: First target of the night, M13 (The Hercules Cluster)… and nothing, because the telescope was off in position by about 50 arcminutes (confirmed by astrometry.net).  As clouds rolled in we did finally get a picture of M13…

And were not able to shutdown calmly:  Maxim DL lost track of its Camera Control Window.  Quitting it and restarting it seemed to make it recover the window.

Thus ends the “interesting” night….

Observing Log, 9/16/2014 0

Posted on September 17, 2014 by Matt

Observed tonight with Nathan Walker, Laura Herzog and Michael Meraz.  The main focus was on reviewing/learning setup and shutdown procedures but we did take some images of M57 (Ring Nebula) in R-band only (to limit time spent on taking flats).

 

Air temperature was about 60F, focus set to 2845 with FWHM of 5.5-6.5.

 

Matt Craig

 



↑ Top